# Intensinist Social Welfare and Ordinal Intensity-Efficient Allocations

### v3

@@ Abstract
We study social welfare and allocation efficiency in situations where, in addition to having ordinal preferences, agents also have “ordinal intensities”: they can make comparisons such as “I prefer $a$ to $b$ more than I prefer $c$ to $d$ without necessarily being able to quantify them. In this new informational environment for social choice we first introduce a rank-based criterion for interpersonal comparisons of such ordinal intensities. Building on it, we introduce the “intensinist” social welfare function. This maps intensity profiles to weak orders over social alternatives using a scoring criterion that generalizes the one in the classic Borda method in a way that allows for differences in agents’ intensities to be reflected in preference aggregation more accurately and in a much richer class of situations. Building on the same comparability postulate we also study the classic assignment problem by defining an allocation to be “intensity-efficient” if it is Pareto efficient with respect to the preferences induced by the agents’ intensities and also such that, when another allocation assigns the same pairs of items to the same pairs of agents but in a “flipped” way, the former allocation assigns the commonly preferred item in every such pair to the agent who prefers it more. We present some first results on the (non-)existence of such allocations without imposing restrictions on preferences or intensities other than strictness, and also study the relation –or lack thereof– between intensity-efficient and classical utilitarian allocations.
@@ 1 Introduction
In Section 5 we again build on the above comparability postulate to study the problem described in our second question. More specifically, assuming here that all agents’ intensity comparisons are strict in the sense that they contain no intensity-equivalances, we define an allocation $x$ as intensity-efficient if it is Pareto efficient and also such that, whenever another allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e. when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in each such pair to the agent who prefers it more. We show that, without any further restrictions, an intensity-efficient allocation exists for all strict intensity profiles when there are three agents and alternatives. Yet with five or more the existence of intensity-efficient allocations is not guaranteed because the underlying dominance relation may be cyclic if additional restrictions are not imposed. We also study the connection –or lack thereof– between intensity-efficient and classical utilitarian allocations in the special informational environment where the latter are well-defined. Perhaps surprisingly, classical utilitarian allocations are not always intensity-efficient. However, we identify an intuitive “balancedness” sufficient condition on utility profiles under which this is indeed the case. An implication of this fact is that, in every situation where the analyst implicitly or explicitly assumes balanced profiles of interpersonally comparable cardinal utility functions whose utility-difference rankings define strict intensity orderings, intensity-efficient allocations always exist.
@@ 5.3 Intensity-Efficient and Utilitarian Allocations Are Distinct
An intensity-efficient allocation exists for any intensity profile that is induced by a interpersonally comparable, strict and balanced cardinal utility profile.
@@
For the second claim, the example strict intensity profile on $A=\{a,b,c,d,e\}$ shown in Table 2 clarifies that an intensity-efficient allocation does not always exist when $n\geq 5$.
@@
Table 2: An intensity profile when $n=5$, for which no intensity-efficient allocation exists.
$s_{i}(a^{\prime},b^{\prime})=$ Agent $i=$ 1 2 3 4 5
10 $(a,e)$ $(a,e)$ $(a,e)$ $(d,\cdot)$ $(e,\cdot)$
9 $(a,d)$ $(a,d)$ $(a,d)$
8 $(b,e)$ $(b,e)$ $(a,c)$
7 $(b,d)$ $(a,c)$ $(b,e)$
6 $(a,c)$ $(b,d)$ $(b,d)$
5 $(c,e)$ $(c,e)$ $(c,e)$
4 $(a,b)$ $(c,d)$ $(c,d)$
3 $(b,c)$ $(a,b)$ $(d,e)$
2 $(c,d)$ $(b,c)$ $(a,b)$
1 $(d,e)$ $(d,e)$ $(b,c)$
In particular, Table 2 shows that
$\begin{array}[]{lllllllll}a&\succ_{i}&b&\succ_{i}&c&\succ_{i}&d&\succ_{i}&e\end{array}$
for $i=1,2,3$, but the intensity orderings of these agents are distinct. On the other hand, the last two columns in Table 2 are meant to be read as suggesting that agents 4 and 5 prefer $d$ and $c$ over everything else, respectively. Other than that, the preferences and intensities of these two agents are inconsequential for the argument. Therefore, Pareto-efficient allocations that correspond to the preferences induced by $(s_{1},\ldots,s_{5})$ are
$\begin{array}[]{lllllll}\underbrace{(c,a,b,d,e)}_{s},&\underbrace{(c,b,a,d,e)}_{t},&\underbrace{(a,b,c,d,e)}_{x},&\underbrace{(a,c,b,d,e)}_{y},&\underbrace{(b,c,a,d,e)}_{w},&\underbrace{(b,a,c,d,e)}_{z}.\end{array}$
[…]
Hence, intensity dominance does not refine the set of Pareto efficient allocations here. $\blacksquare$
The somewhat surprising fact that a plausible dominance concept may by cyclic and prevent an optimal entity to emerge invites an informal analogy to be drawn between intensity-dominance cycles over allocations with at least 5 agents and Condorcet cycles over alternatives in majority-based preference aggregation with at least 3 agents (Sen, 2017/1970). Unlike that framework, however, although intensity-dominance cycles here may prevent refining the Pareto set, they do not lead to a “policy paralysis” problem because Pareto allocations always exist and, absent any distributionally juster suggestions, one of them might be promoted by the social planner.
5.5 Example ($n=4$): Intensity-Efficiency Refines Pareto

### v4

@@ Abstract
This paper studies social welfare and allocation efficiency in situations where, in addition to having ordinal preferences, agents also have ordinal intensities: they can make comparisons such as “I prefer $a$ to $b$ more than I prefer $c$ to $d$ without necessarily being able to quantify them. In this new informational environment for social choice, the paper first introduces a rank-based criterion for interpersonal comparisons of such ordinal intensities. Building on it, the “intensinist” social welfare function is introduced. This maps profiles of ordinal intensities to weak orders over social alternatives using a scoring method which generalizes that of the classic Borda count in a way that allows for differences in agents’ intensities to be reflected in preference aggregation more accurately and in a much richer class of situations. Building on the same comparability criterion, the paper also studies the classic assignment problem by defining an allocation to be “intensity-efficient” if it is Pareto efficient with respect to the preferences induced by the agents’ intensities and also such that, when another allocation assigns the same pairs of items to the same pairs of agents but in a “flipped” way, the former allocation assigns the commonly preferred item in every such pair to the agent who prefers it more. Some first results on the (non-)existence of such allocations are presented without imposing restrictions on preferences or intensities other than strictness, and the relation –or lack thereof– is studied between intensity-efficient and classical utilitarian allocations.
@@ 1 Introduction
In Section 5 we again build on the above comparability postulate to study the problem described in our second question. More specifically, assuming here that all agents’ intensity comparisons are strict in the sense that they contain no intensity-equivalances, we define an allocation $x$ as intensity-efficient if it is Pareto efficient and also such that, whenever another allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e. when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in each such pair to the agent who prefers it more. We show that, without any further restrictions, an intensity-efficient allocation exists for all strict intensity profiles when there are three agents and alternatives. Yet we also show that with four or more the existence of intensity-efficient allocations is not guaranteed because the underlying dominance relation may be cyclic if additional restrictions are not imposed. We also study the connection –or lack thereof– between intensity-efficient and classical utilitarian allocations in the special informational environment where the latter are well-defined. Perhaps surprisingly, classical utilitarian allocations are not always intensity-efficient. However, we identify an intuitive “balancedness” sufficient condition on utility profiles under which this is indeed the case. An implication of this fact is that, in every situation where the analyst implicitly or explicitly assumes balanced profiles of interpersonally comparable cardinal utility functions whose utility-difference rankings define strict intensity orderings, intensity-efficient allocations always exist.
@@ 5.3 Intensity-Efficient and Utilitarian Allocations Are Distinct
An intensity-efficient allocation exists for any intensity profile that is induced by an interpersonally comparable, strict and balanced cardinal utility profile.
@@
For the claimed potential non-existence when $n\geq 4$, consider the example strict intensity profile on $A=\{a,b,c,d\}$ whose canonical representation is shown in the table below:
@@
6 $(a,d)$ $(a,d)$ $(a,d)$ $(a,d)$
5 $(b,d)$ $(a,c)$ $(a,c)$ $(b,d)$
4 $(a,c)$ $(b,d)$ $(b,d)$ $(a,c)$
3 $(b,c)$ $(b,c)$ $(b,c)$ $(b,c)$
2 $(a,b)$ $(c,d)$ $(c,d)$ $(a,b)$
1 $(c,d)$ $(a,b)$ $(a,b)$ $(c,d)$
We start by observing that $s_{1}=s_{4}$, $s_{2}=s_{3}$ and
$\begin{array}[]{ccccccccc}a&\succ_{i}&b&\succ_{i}&c&\succ_{i}&d,&&i=1,\ldots,4\end{array}$
This implies that all 24 possible allocations, defined and listed below, are Pareto efficient.
$\begin{array}[]{llllllllllllllllll}x_{1}&=&(a,b,c,d)&&&x_{2}&=&(a,b,d,c)&&&x_{3}&=&(a,c,b,d)&&&x_{4}&=&(a,c,d,b)\\ x_{5}&=&(a,d,b,c)&&&x_{6}&=&(a,d,c,b)&&&x_{7}&=&(b,a,c,d)&&&x_{8}&=&(b,a,d,c)\\ x_{9}&=&(b,c,a,d)&&&x_{10}&=&(b,c,d,a)&&&x_{11}&=&(b,d,a,c)&&&x_{12}&=&(b,d,c,a)\\ x_{13}&=&(c,a,b,d)&&&x_{14}&=&(c,a,d,b)&&&x_{15}&=&(c,b,a,d)&&&x_{16}&=&(c,b,d,a)\\ x_{17}&=&(c,d,a,b)&&&x_{18}&=&(c,d,b,a)&&&x_{19}&=&(d,a,b,c)&&&x_{20}&=&(d,a,c,b)\\ x_{21}&=&(d,b,a,c)&&&x_{22}&=&(d,b,c,a)&&&x_{23}&=&(d,c,a,b)&&&x_{24}&=&(d,c,b,a)\end{array}$
The following comparisons, whose validity can be readily established by the reader, demonstrate that for each $x_{i}$, $i=1,\ldots,24$, in this set there is a distinct $x_{j}$ that intensity-dominates $x_{i}$ (a situation denoted by $x_{j}Dx_{i}$).
$\begin{array}[]{llllllll}x_{6}Dx_{1}&x_{1}Dx_{2}&x_{4}Dx_{3}&x_{13}Dx_{4}&x_{3}Dx_{5}&x_{17}Dx_{6}&x_{1}Dx_{7}&x_{2}Dx_{8}\\ x_{3}Dx_{9}&x_{8}Dx_{10}&x_{9}Dx_{11}&x_{11}Dx_{12}&x_{18}Dx_{13}&x_{13}Dx_{14}&x_{17}Dx_{15}&x_{22}Dx_{16}\\ x_{23}Dx_{17}&x_{24}Dx_{18}&x_{8}Dx_{19}&x_{22}Dx_{20}&x_{11}Dx_{23}&x_{12}Dx_{22}&x_{24}Dx_{23}&x_{10}Dx_{24}\end{array}$
Therefore, no intensity-efficient allocation exists in this profile. $\blacksquare$
The somewhat surprising fact that a plausible dominance concept may by cyclic and prevent an optimal entity to emerge invites an informal analogy to be drawn between intensity-dominance cycles over allocations with at least 4
agents and Condorcet cycles over alternatives in majority-based preference aggregation with at least 3 agents (Sen, 2017/1970). Unlike that framework, however, although intensity-dominance cycles here may prevent refining the Pareto set, they do not lead to a “policy paralysis” problem because Pareto allocations always exist and, absent any distributionally juster suggestions, one of them might be promoted by the social planner.
5.5 Example where Intensity-Efficiency Refines Pareto ($n=4$)